Circular No. 7828

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2001 YX_{127} (LINEAR) An apparently asteroidal object discovered by LINEAR (discovery ob-

servation below) and designated 2001 YX_{127} (cf. MPS 47220, MPO 24028) has been found to have a broad, fan-shaped extension in p.a. 100^{o} on CCD images obtained on Feb. 14.2 UT by T. B. Spahr with the 1.2-m reflector at Mount Hopkins. Co-added CCD R-band images taken at about the same time by C. W. Hergenrother with the Catalina 1.54-m reflector show a 7′′ coma and a broad tail 8′′ long in p.a. 100^{o}.

2001 UT  α_{2000}  δ_{2000}  m_{2}
Dec. 17.31658  4^{h}47^{m}24.^{s}54  +30^{o}02′59′′.7  20.2

Complete astrometry, the following orbital elements, and an ephemeris appear on MPEC 2002-C97:

T = 2003 Mar. 13.2865 TT  ω = 116.^{o}5924
e = 0.179914  Ω = 31.5250 }2000.0
q = 3.419865 AU  i = 7.9131
a = 4.170129 AU n^{o} = 0.1157390 P = 8.516 years

SUPERNOVA 2002an IN NGC 2575 S. Benetti, G. Altavilla, A. Pastorello, M. Turatto, S. Desidera, and E.

Giro, Osservatorio Astronomico di Padova; and E. Cappellaro, Osservatorio Astronomico di Capodimonte, write: “A spectrum (range 380–780 nm, resolution 2.5 nm) of SN 2002an (cf. IAUC 7805), obtained on Feb. 5.02 UT with the Asiago 1.82-m telescope (+ AFOSC), shows that it is a normal type-II supernova about a week past maximum. The spectrum is dominated by a blue continuum: spectral model fitting using SYNOW (Jeffery and Branch 1990, in Supernovae, World Scientific, p. 149) gives a blackbody temperature of ∼ 11 000 K and expansion velocity of ∼ 9 000 km/s, as deduced from H I, He I, and Ba II P-Cyg line profiles. The P-Cyg profile of Hα is not yet well developed, mostly consisting of a broad shallow emission. Evidence of interaction with the circumstellar medium (cf. IAUC 7818) is not visible at this epoch.”

V2540 OPHIUCHI Visual magnitude estimates by A. Pearce, Nedlands, W. Australia:

Jan. 30.842 UT, 8.4; Feb. 7.833, 9.8; 10.856, 9.4; 11.826, 8.2; 13.837, 8.5.

2002 February 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green